# Meaningful Learning: Advancing Abstract Reasoning in Large Language Models via Generic Fact Guidance

### v1

@@ 5 Experiments
5.1 Training Details
We use LoRA Hu et al. (2022) for parameter effcient finetuning of Orca-2. The batch size is set as 256 and 240 for 7B and 13B Orca-2, respectively. We train them for one epoch with a learning rate of 5e-5. 7B and 13B Orca-2 are respectively trained on 2 and 3 NVIDIA A100 80GB PCIe GPUs.
5.2 Evaluation Benchmarks
To evaluate the MeanLearn paradigm, in addition to AbsR, we also include a wide range of benchmarks of reasoning and natural language understanding:
[…]
5.3 Evalution Settings
Since the generation-based evaluation is time-consuming due to the limitations of computing resources, we follow OpenCompass Contributors (2023) to have hybrid evaluation criteria:
(1) For free-form generation tasks (4 tasks in BBH), we ask LLMs to generate no more than 250 tokens with a temperature of 0.0, and then we utilize exact match for evaluation.
(2) For the other classification and mutipul-choice questions, perplexity (PPL) of each category or option is adopted for prediction. The evaluation templates can refer to Appendix E. Take Orca-2 and MeanLearn as an example, we use:
<|im_start|> system{INSTRUCTION} <|im_end|> <|im_start|> userQuestion:{QUESTION} <|im_end|> <|im_start|> assistantAnswer:{OPTION} <|im_end|>
The text enclosed within curly braces is a placeholder. {INSTRUCTION} will be filled with task-specific instruction. {QUESTION} will be filled with the question and all candidate options or categories. {OPTION} will be filled with the option index only (e.g. A).
To measure the performance, the option or category with the lowest PPL will be chosen as the prediction of corresponding LLM. The performance of each benchmark is weighted across all tasks of datasets in the benchmark.
Size LLMs AbsR AGIEval RACE BBH Commonsense MMLU ARC-e ARC-c Average
@@ 5.3 Evalution Settings
7B LLaMA-2 50.00 27.89 38.85 26.00 55.37 41.65 58.73 43.05 42.69
Orca-2 73.50 38.57 73.32 34.51 71.58 50.11 79.19 73.90 61.84
MeanLearn 77.00 38.33 77.22 35.71 76.55 53.04 86.60 78.31 65.35
13B LLaMA-2 73.50 34.41 59.46 30.61 61.00 51.87 71.25 55.25 54.67
Orca-2 66.50 43.75 66.86 39.39 65.92 47.27 82.19 70.85 60.34
MeanLearn 66.00 43.95 70.47 40.45 69.93 51.62 89.59 82.03 64.26
@@
7B LLaMA-2 35.42 16.09 14.08 9.56 47.42 25.48 50.27 35.51 29.23
Orca-2 60.42 26.27 50.90 15.67 64.02 33.68 72.16 67.76 48.86
MeanLearn 64.58 25.27 57.10 16.67 71.17 37.24 81.35 73.83 53.40
13B LLaMA-2 61.46 21.01 32.17 13.74 52.87 35.72 62.43 44.86 40.53
Orca-2 50.00 29.50 42.12 20.61 56.41 31.11 75.95 62.15 45.98
MeanLearn 45.83 30.06 47.89 22.00 61.06 35.24 85.41 77.10 50.57
Table 6: The overall vanilla accuracy and AbsAcc (%) of all LLMs.
@@ 5.4 Results: Vanilla Accuracy
Table 6 shows the Vanilla Accuracy of MeanLearn and baselines on AbsR and the benchmarks. We can infer that:
(1) Orca-2 siginificantly outperforms LLaMA-2 on most benchmarks, which indicates that massive reasoning-specific data can teach LLMs to develop much better reasoning abilities.
(2) MeanLearn has superiority over LLaMA-2 and Orca-2 on nearly all benchmarks. This demonstrates MeanLearn can use generic facts to effectively guide LLMs to reason more properly and logically, resulting in better reasoning capabilities.
(3) The more challenging the benchmark (e.g. BBH), the less pronounced the improvement in the performance of MeanLearn. This is mainly because the improvement is mostly constrained by the complex scenarios in harder benchmarks, which requires mightier base LLMs or more complex training data to perform and learn multi-step and compositional reasoning.
(4) Post-training (Orca-2 and MeanLearn) does not always bring improvements (e.g. AGIEval). The primary reason is the reasoning-specific post-training, which is bound to affect the general abilities and the handle of uncovered reasoning types.
(5) Larger LLMs do not always yield better results. We suppose the main factors behind this are in two ways: on the one hand, LLMs with different parameter sizes might be good at different tasks, excessive contemplation may lead to decreases on some tasks Wei et al. (2022b). On the other hand, the perplexity-based evaluation setting might has moderate cross-scale generalization ability.
@@ 5.5 Results: Abstract Reasoning Accuracy
Since the generic facts are not given in the benchmarks, we train a RoBERTa-Large Liu et al. (2019) with constrastive learning to cluster examples supported by the same generic fact. In particular, we utilize e-CARE as the training set, examples guided by a same generic fact will be pushed closer, or they will be pushed away. Cosine similarity is used to measure the distance between examples. For clustering, examples sharing a similarity above 0.6 are considered to be supported by the same generic fact. Each cluster contains no more than 3 examples.
@@
(1) The AbsAcc is much lower than the vanilla accuracy, indicating the large gap between general reasoning and abatract reasoning. Perhaps more training data can help to fill the gap between vanilla accuracy and abstract reasoning accuracy.
(2) MeanLearn surpasses Orca-2 and LLaMA-2, exhibiting greater advantages in AbsAcc than in vanilla accuracy. This suggests that the improvement in vanilla accuracy is more likely due to the enhancement of abstract reasoning abilities. MeanLearn can implicitly tell LLMs the generic fact and teach them how to utilize it, encouraging LLMs focus more on abstract reasoning.
6 Analysis
To further investigate the pros, cons, and effectiveness of MeanLearn and baselines, we conduct several in-depth analyses and ablation studies.
Size LLMs AbsR AGIEval RACE BBH Commonsense MMLU ARC-e ARC-c Average
@@ 6.1 Performance on Different Domains
For potential future improvements, we categorize tasks of MMLU based on their respective domains to explore the domains where MeanLearn and baselines excel and where they does not excel. The results are shown in Figure 4 with a total of 16 categories (e.g. law), from which we can conclude:
@@
6.2 Effect of Knowledge and Reasoning
@@
(4) In AGIEval, Knowledge and Reasoning are less critical due to its focus on complex reading comprehension without needing vast knowledge but. Moreover, Knowledge seems useless in Commonsense. When providing LLMs with knowledge, we should also teach them how to use it.
6.3 Effect of Guided Explanations
When constructing AbsR, we would ask GPT-4 to generate an explanation guided by the generic fact (Sec 3.2). Hence, we devise another ablation study to investigate the effect of the guidance brought by the generic fact.
To be precise, we follow the process of generating samples with GPT-4 in Sec 3.2, but we do not provide the generic fact in the prompt. Therefore, the explanation is guided by GPT-4 itself. We denote this dataset as $*{}^{*}$. Results are shown in Table 7 (w/o Guidance), from which we can infer:
(1) Training 7B and 13B Orca-2 with $*{}^{*}$ can still outperform Orca-2 on vanilla accuracy. However, it cannot bring stable improvements on AbsAcc. This underscores that reasoning-specific post-training can boost performance, while improving abstract reasoning performance might depend on the difficulty of the benchmarks.
(2) MeanLearn without Knowledge outperforms MeanLearn without Guidance, showing the superiority of training with generic-fact-guided explanation. This stems from the generic-fact-guided explanation that can teach LLMs to use the regular pattern behind the questions for better reasoning.
@@ 7.1 Reasoning with LLMs
As for prompt-based methods, Wei et al. (2022b) introduced CoT prompt to elicit reasoning in LLMs by encouraging step-by-step thinking. Additionally, Kojima et al. (2022) presented zero-shot CoT to bypass manual annotation issues. Subsequently, Zhang et al. (2023) proposed AutoCoT to automatically generate few-shot examples with CoT for reasoning. Various CoT-based methods have since emerged, such as complex CoT Fu et al. (2023), and tree of thought Yao et al. (2023), etc.
@@ 7.2 LLMs-as-Annotators
LLMs are increasingly employed for data labeling due to their advanced ability to follow instructions.
Some studies distilled various CoT data from LLMs to train student models (Li et al., 2023; Ho et al., 2023; inter alia), achieving impressive performance. For instance, Li et al. (2023) distilled symbolic CoT from GPT-3 and observed enhanced performance in commonsense reasoning. With the advent of open-source LLMs like LLaMA Touvron et al. (2023), other studies distilled instruction data from large-scale LLMs to train small-scale LLMs (Xu et al., 2023; Mukherjee et al., 2023; inter alia). For example, Xu et al. (2023) employed an evolving strategy to obtain complex data, training LLaMA and achieving strong performance.
We employ LLMs to generate training data and develop AbsR tailored for abstract reasoning.
@@ 8 Conclusion
In this paper, we investigate the abstract reasoning of open-source LLMs. To achieve this, we first define the evaluation metric of abstract reasoning and design a series of preliminary experiments. Then we discover a large performance gap between general reasoning and abstract reasoning. Hence, we tailor an abstract reasoning dataset AbsR with the help of GPT-4 to enhance LLMs. Finally, we devise a simple but effective paradigm (MeanLearn) to teach LLMs abstract reasoning in Knowledge and Reasoning ways. Extensive experiments demonstrate the superiority and effectiveness of MeanLearn in vanilla and abstract reasoning accuracies.
Limitations
There are several limitations of our work: First, the size of the training data should be larger to obtain better and more stable results. Second, since the large performance improvement might be due to the PPL-based evaluation, the evaluation setting should be in a generative way to get more precise results. There should be some examples (whether MeanLearn can generate meaningful explanations) to demonstrate the superiority of MeanLearn. Third, more efforts should be made to focus more on abstract reasoning itself. Finally, more LLMs should be incorporated for training to demonstrate the generalizability of MeanLearn.
Appendix A Asbstract Reasoning Study
A.1 Prompt for Experiment I
<|im_start|> system You are Orca, an AI language model created by Microsoft. You are a cautious assistant. You carefully follow instructions. You are helpful and harmless and you follow ethical guidelines and promote positive behavior.<|im_end|> <|im_start|> user{premise} Hypothesis1 or Hypothesis2? Hypothesis1: {hypothesis1} Hypothesis2: {hypothesis2} Your answer should follow the format like ‘‘Answer: Hypothesis(1 or 2) is more plausible. Explanation: ___’’<|im_end|> <|im_start|> assistant

### v2

@@ 5 Experiments
5.1 Experimental details
Training setup. We use LoRA [20] for parameter effcient finetuning. 7B and 13B MeanLearn are trained on 7B and 13B Orca-2, respectively, 8B MeanLearn is trained on 8B LLaMA-3, with batch sizes of 256 for 7B and 8B, and 240 for 13B. MeanLearn of various sizes are trained for one epoch at a 5e-5 learning rate with AdamW [26]. 7B and 8B MeanLearn are trained on 2 NVIDIA A100 80GB PCIe GPUs for 3 hours. 13B MearnLearn is trained on 3 such GPUs for 4 hours.
Baselines. We adopt a wide range of open and limited access LLMs across different sizes as baselines. Open Access LLMs: (1) LLaMA-2 [47] 7B and 13B; (2) LLaMA-3 [34] 8B; (3) Vicuna [8] 7B and 13B (finetuned on ShareGPT [1] data); (4) WizardLM [56] 7B and 13B (fintuned with evolved instruction data); (5) Orca-2 [35] 7B and 13B (prograssively finetuned on massive reasoning data). Limited Access LLMs: (1) GPT-3.5 (gpt-3.5-turbo-0125); (2) PaLM-2 (text-bison-001).
Evaluation benchmarks. In addition to AbsR, we also include a wide range of benchmarks of reasoning and natural language understanding for evaluation: (1) AGIEval [60] consists of tests ranging from college admission tests, to national civil service examinations; (2) RACE [28] consists of tests with reading comprehension questions; (3) BBH [43] is a subset of Big-Bench [42], which contains 23 hardest tasks focusing on challenging scenarios; (4) Com. [55] is a collection of 7 commonsense reasoning datasets ($\alpha$NLI [4], CSQA [44], COPA [39], e-CARE [14], SocialIQa [40], PIQA [6], and StrategyQA [16]); (5) MMLU [18] is a massive multitask language understanding benchmark; (6) ARC [10] is a benchmark of easy (ARC-e) and challenge (ARC-c) science questions.
Evaluation setup. Since generation-based evaluation is time-consuming due to limited computing resources, we follow OpenCompass [11] to have hybrid evaluation criteria with high reproducibility: (1) For PaLM-2 and GPT-3.5 on all tasks, and open access LLMs on 4 generation tasks (4 tasks in BBH), greedy decoding and exact match are utilized for evaluation; (2) For open access LLMs on classification and multiple-choice tasks, perplexity (PPL) is adopted for prediction. The option or category with the lowest PPL is chosen as the answer. For all baselines, evaluation is conducted once since there is no randomness on such evaluation criteria. For MeanLearn, results are averaged of MeanLearn(s) trained with 3 different random seeds. The evaluation prompts can refer to Appendix H.
Table 5: Overall vanilla accuracy and AbsAcc (%) of baselines and MeanLearn. Due to the space limit, we only report the standard deviation of Average performance of each method.
Size LLMs AbsR AGIEval RACE BBH Com. MMLU ARC-e ARC-c Average
@@ 5.1 Experimental details
>20B PaLM-2 75.76 15.83 75.82 48.73 78.05 58.73 89.95 82.37 65.77 $\pm$ 0.00
GPT-3.5 84.00 25.84 83.36 54.15 74.80 65.98 94.71 88.81 71.46 $\pm$ 0.00
7B LLaMA-2 50.00 27.89 38.85 26.00 55.37 41.65 58.73 43.05 42.69 $\pm$ 0.00
Vicuna 75.00 32.56 65.75 31.93 64.77 49.40 74.78 57.63 56.48 $\pm$ 0.00
WizardLM 65.00 22.27 22.11 25.99 43.63 23.26 25.57 22.37 31.28 $\pm$ 0.00
Orca-2 73.50 38.57 73.32 34.51 71.58 50.11 79.19 73.90 61.84 $\pm$ 0.00
MeanLearn 77.00 38.15 77.15 35.64 76.59 52.98 86.67 78.06 65.28 $\pm$ 0.34
8B LLaMA-3 82.50 36.33 76.75 36.60 71.76 61.86 88.54 75.25 66.20 $\pm$ 0.00
MeanLearn 84.50 38.53 77.02 37.67 71.73 62.35 88.89 77.97 67.12 $\pm$ 0.18
13B LLaMA-2 73.50 34.41 59.46 30.61 61.00 51.87 71.25 55.25 54.67 $\pm$ 0.00
Vicuna 74.00 33.23 63.54 33.22 63.10 49.34 77.78 59.77 56.75 $\pm$ 0.00
WizardLM 74.00 33.23 63.54 33.22 63.10 49.34 77.78 59.77 56.75 $\pm$ 0.00
Orca-2 66.50 43.75 66.86 39.39 65.92 47.27 82.19 70.85 60.34 $\pm$ 0.00
MeanLearn 67.17 43.92 70.24 40.36 69.75 51.68 88.69 82.00 64.23 $\pm$ 0.25
@@
>20B PaLM-2 64.58 9.59 61.54 27.69 69.34 44.39 85.68 77.10 54.99 $\pm$ 0.00
GPT-3.5 77.08 14.48 66.79 25.45 65.74 52.09 92.16 85.05 59.86 $\pm$ 0.00
7B LLaMA-2 35.42 16.09 14.08 9.56 47.42 25.48 50.27 35.51 29.23 $\pm$ 0.00
Vicuna 58.33 19.90 38.68 14.28 58.07 32.50 65.95 49.53 39.84 $\pm$ 0.00
WizardLM 53.24 12.94 5.57 9.77 32.96 11.55 20.00 17.29 20.42 $\pm$ 0.00
Orca-2 60.42 26.27 50.90 15.67 64.02 33.68 72.16 67.76 48.86 $\pm$ 0.00
MeanLearn 64.58 25.27 57.10 16.67 71.17 37.24 81.35 73.83 53.39 $\pm$ 0.32
8B LLaMA-3 72.92 23.78 55.61 18.13 66.35 47.31 83.51 70.09 54.71 $\pm$ 0.00
MeanLearn 73.96 26.57 55.09 19.24 66.48 46.76 83.78 73.36 55.66 $\pm$ 0.17
13B LLaMA-2 61.46 21.01 32.17 13.74 52.87 35.72 62.43 44.86 40.53 $\pm$ 0.00
Vicuna 61.46 19.70 36.50 16.33 55.26 32.21 70.27 52.34 40.37 $\pm$ 0.00
WizardLM 59.38 16.96 28.59 18.61 55.73 29.47 57.84 42.06 38.58 $\pm$ 0.00
Orca-2 50.00 29.50 42.12 20.61 56.41 31.11 75.95 62.15 45.98 $\pm$ 0.00
MeanLearn 45.82 30.01 47.65 21.85 61.04 35.21 85.41 77.08 50.51 $\pm$ 0.23
@@ 5.2 Results: vanilla accuracy
Table 5 shows the Vanilla Accuracy of MeanLearn and baselines:
(1) Additional training usually brings improvements  (Vicuna, WizardLM, Orca-2, and MeanLearn), but it still can result in performance losses on some benchmarks (e.g. AGIEval). This signifies the effectiveness of massive reasoning-specific data and the necessity of data coverage.
(2) MeanLearn has superiority over all open access baselines on nearly all benchmarks. This demonstrates that MeanLearn can use generic facts to effectively guide LLMs to reason more properly and logically, making them more flexible and task-adaptive.
(3) The more challenging the benchmark (e.g. BBH), the less pronounced the improvement in the performance of MeanLearn. This is mainly because the improvement is mostly constrained by the complex scenarios in harder benchmarks, which require mightier base LLMs or more complex training data to perform and learn multi-step and compositional reasoning.
(4) MeanLearn does not have many advantages over LLaMA-3. LLaMA-3 stores dense knowledge in its parameters, making it harder to learn new knowledge without forgetting existing knowledge.
(5) Larger open-sourced LLMs do not always yield better results. We suppose the main factors behind this are in two ways: on the one hand, LLMs with different parameter sizes might be good at different tasks, and excessive contemplation may lead to decreases on some tasks [53]. On the other hand, the perplexity-based evaluation setting might have moderate cross-scale generalization ability.
(6) LLMs possess more than 20B parameters still struggle with abstract reasoning. MeanLearn can achieve comparable performance to PaLM-2 despite their huge gap in model sizes.
@@ 5.3 Results: abstract reasoning accuracy
Since the generic facts are not given in the OOD benchmarks, we train a RoBERTa-Large [32] with constrastive learning to cluster examples supported by the same generic fact. In particular, we utilize e-CARE as the training set, examples guided by the same generic fact will be pushed closer, or they will be pushed away. Cosine similarity is used to measure the distance between examples (training details can refer to Appendix I). For clustering, examples sharing a similarity above 0.6 are considered to be supported by the same generic fact. Each cluster contains no more than 3 examples.
@@
(1) AbsAcc is much lower than vanilla accuracy, indicating the large gap between general reasoning and abstract reasoning. More efforts should be made to fill this gap in the future.
(2) MeanLearn can surpass all baselines, usually exhibiting greater advantages in AbsAcc than in vanilla accuracy. This suggests that the improvement in vanilla accuracy is more likely due to the enhancement of abstract reasoning abilities. MeanLearn can tell LLMs the generic facts and teach them how to implicitly utilize them in different scenarios, resulting in better abstract reasoning.
6 Further analysis
To further investigate the pros, cons, and effectiveness of MeanLearn, we conduct several analyses and ablation studies. LLaMA-2, Orca-2, and MeanLearn are adopted for analysis.
Size Method AbsR AGIEval RACE BBH Com. MMLU ARC-e ARC-c Average
@@ 6.1 Performance on different domains
We categorize tasks of MMLU based on their respective domains to demonstrate the pros and cons of MeanLearn for future improvements. The results are shown in Figure 4 with a total of 16 categories:
@@
6.2 How much do knowledge and reasoning ways contribute to MeanLearn?
@@
(4) In AGIEval, Knowledge and Reasoning are less critical due to their focus on complex reading comprehension without needing vast knowledge. Moreover, Knowledge seems useless in Commonsense. When providing LLMs with knowledge, we should also teach them how to use it.
6.3 Is the improvement in performance solely attributed to additional data?
Supervised fine-tuning can also improve AbsAcc (Table 1), so we design another ablation study to investigate whether the improvement of MeanLearn only comes from additional data.
To be precise, for each sample in AbsR, we use GPT-4 to generate a new explanation based on the question and label without the guidance of the generic fact. Hereafter, we obtain AbsR. The only difference between AbsR and AbsR is the explanations in AbsR are guided by the generic fact. Table 6 (w/ AbsR) shows the results of MeanLearn trained with AbsR, we can infer:
(1) Training 7B and 13B Orca-2 with AbsR can still outperform Orca-2 on vanilla accuracy. However, it cannot bring stable improvements on AbsAcc. This underscores that reasoning-specific post-training can boost performance, while improving abstract reasoning performance might depend on more careful design such as MeanLearn.
(2) MeanLearn without Knowledge outperforms MeanLearn trained with AbsR, showing the superiority of training with generic-fact-guided explanation. This stems from the generic-fact-guided explanation that can teach LLMs to use the regular pattern behind the questions for better reasoning.
Table 7: MeanLearn trained based on Mistral.
Method AbsR Com. MMLU RACE Average
Vanilla Accuracy
Mistral 83.00 66.53 57.02 75.33 70.47
MeanLearn (Ours) 84.50 74.56 58.28 76.65 73.50
AbsAcc
Mistral 72.92 58.87 40.19 52.69 56.17
MeanLearn (Ours) 73.96 68.59 41.14 55.77 59.87
Table 8: Performance on mathematical reasonin datasets from MMLU [18].
Size Method AA CM EM HSS HSM Average
Vanilla Accuracy
7B Orca-2 23.00 32.00 33.07 37.95 28.15 30.83
MeanLearn 36.00 35.00 33.07 40.28 28.89 34.56
Mistral 25.00 30.00 38.89 46.76 32.96 34.72
MeanLearn 31.00 34.00 37.83 47.69 34.07 36.92
AbsAcc
7B Orca-2 0.00 4.88 25.00 25.17 14.63 13.94
MeanLearn 2.86 7.32 24.02 27.97 16.26 15.69
Mistral 2.86 7.32 25.49 36.36 14.63 17.33
MeanLearn 5.71 9.76 24.55 37.76 16.26 18.81
6.4 Expanding MeanLearn to LLMs besides the LLaMA series
To investigate the universality of MeanLearn, we train MeanLearn with Mistral-7B-Instruct-v0.2 [21] (denoted as Mistral). All settings are consistent with Sec 5. We choose AbsR, Com., MMLU, and RACE for evaluation. As shown in Table 8, we can find MeanLearn has good applicability on Mistral.
6.5 Expanding reasoning domains to mathematical reasoning
Although we do not incorporate math examples in AbsR, we want to investigate whether MeanLearn can improve the performance of mathematical reasoning by enhancing abstract reasoning in LLMs. Following [33], we select five mathematical reasoning datasets from MMLU [18] to demonstrate this: abstract algebra (AA), college mathematics (CM), elementary mathematics (EM), high school statistics (HSS), and high school mathematics (HSM). We choose 7B version of Orca-2 and Mistral for comparison. Results are shown in Table 8, it is interesting to find that MeanLearn can outperform baselines even if there is no math data in AbsR. This demonstrates the effectiveness of MeanLearn in enhancing abstract reasoning in LLMs.
@@ 7.1 Reasoning with LLMs
As for prompt-based methods, Wei et al. [53] introduced CoT prompt to elicit reasoning in LLMs by encouraging step-by-step thinking. Additionally, Kojima et al. [27] presented zero-shot CoT to bypass manual annotation issues. Subsequently, Zhang et al. [59] proposed AutoCoT to automatically generate few-shot examples with CoT for reasoning. Various CoT-based methods [22, 51] have since emerged, such as complex CoT [15], and the tree of thought [57], etc.
@@ 7.2 LLMs-as-annotators
LLMs are increasingly used as annotators to create data for training and evaluation due to their advanced instruction-following skills.
Some studies distilled various CoT data from LLMs to train student models ([30, 19, 58], inter alia), achieving impressive performance. For instance, Li et al. [30] distilled symbolic CoT from GPT-3 and observed enhanced performance in commonsense reasoning. Dai et al. [12] used ChatGPT to generate augmentation data to enhance BERT [24]. With the advent of open-source LLMs like LLaMA [47], other studies distilled instruction data from large-scale LLMs to train small-scale models ([56, 36, 25, 31, 23, 17], inter alia). For example, Xu et al. [56] and Luo et al. [33] employed an evolving strategy to obtain complex data, trained LLaMA, and achieved strong performance. Tang et al. [45] constructed a concept graph and used GPT-3.5 to synthesize math data with high diversity. We employ LLMs to generate training data and develop AbsR tailored for abstract reasoning.
@@ 8 Conclusion
In this paper, we investigate the abstract reasoning of existing LLMs. To achieve this, we first define the evaluation metric of abstract reasoning and design a series of preliminary experiments. Then we discover a significant performance gap between general reasoning and abstract reasoning. Hence, we tailor an abstract reasoning dataset AbsR with the help of GPT-4 to enhance LLMs. Finally, we devise a simple but effective paradigm (MeanLearn) to teach LLMs abstract reasoning in Knowledge and Reasoning ways. Extensive experiments demonstrate the superiority and effectiveness of MeanLearn in vanilla and abstract reasoning accuracies. The limitations of MeanLearn are discussed in Appendix A.
Acknowledgements
We would like to acknowledge the editors and reviewers for their efforts and advice, and gratefully acknowledge the support of the National Natural Science Foundation of China under Grants U22B2059 and 62176079, Natural Science Foundation of Heilongjiang Province under Grant YQ2022F005.
Appendix A Limitations
There are several limitations of our work: First, the size of the training data should be larger to obtain better and more stable results. Second, there should be some examples (whether MeanLearn can generate meaningful explanations) to demonstrate the superiority of MeanLearn. Third, more efforts should be made to focus more on abstract reasoning itself. Finally, complex scenarios (with complex generic facts) such as causal chain reasoning [54] and logical reasoning [41] is a potential direction of abstract reasoning.
Appendix B Asbstract reasoning study
B.1 Prompt for experiment I
B.1.1 Orca-2
<|im_start|> system
You are Orca, an AI language model created by Microsoft. You are a cautious assistant. You carefully follow instructions. You are helpful and harmless and you follow ethical guidelines and promote positive behavior.<|im_end|>
<|im_start|> user
{premise} Hypothesis1 or Hypothesis2?
Hypothesis1: {hypothesis1}
Hypothesis2: {hypothesis2}
Your answer should follow the format like ‘‘Answer: Hypothesis(1 or 2) is more plausible.
Explanation: ___’’<|im_end|>
<|im_start|> assistant
B.1.2 LLaMA-2 and GPT-3.5
System : You are a helpful assistant.
User: {premise} Hypothesis1 or Hypothesis2?
Hypothesis1: {hypothesis1}
Hypothesis2: {hypothesis2}
Your answer should follow the format like ‘‘Answer: Hypothesis(1 or 2) is more plausible.
Explanation: ___’’